OpenStudy (anonymous):

Which is a solution of 3(x + 4) − x > 4?

OpenStudy (anonymous):

x>0

OpenStudy (anonymous):

?

OpenStudy (anonymous):

A. –6 B. –5 C. –4 D. –3

OpenStudy (anonymous):

so, u distribute the 3

OpenStudy (anonymous):

my bad 4

OpenStudy (anonymous):

C is the answer

OpenStudy (kl0723):

3x+12-x>4 3x-x>4-12 2x=-8 x=-8/2 x=-4

OpenStudy (anonymous):

the ?s are -

OpenStudy (kl0723):

I threw an equal sign lol is 4 then ;P

OpenStudy (anonymous):

it is actrually -4, but according to ur context, i'd say 4